OpenStudy (anonymous):

solve for t: -pi/4 * sin(pi*t/4) = 0

OpenStudy (anonymous):

-pi/4 * sin(pi*t/4) = 0 sin(pi*t/4) = 0 pi*t/4 = 0 or pi or 2pi or ...... t=0 or 4 or 8 or....